OpenStudy (anonymous):

Simplify the following expression. √-6√-42

OpenStudy (anonymous):

Start by getting rid of the minus signs like this sqrt ( -6) = sqrt(-1)sqrt(6) = i sqrt 6

OpenStudy (anonymous):

sqrt(-43)= sqrt(-1)sqrt=isqrt43

OpenStudy (anonymous):

sqrt(6)i * sqrt(42)i get values of sqrt(6) and sqrt(42) from calculator.. or calculate yourself for self evaluation :D

OpenStudy (anonymous):

42, right?

OpenStudy (anonymous):

Now the i*i u deal with before anything else, i^2 = i*i = -1

OpenStudy (anonymous):

so sqrt(6)i*(42)i

OpenStudy (anonymous):

a calculator will not help you hear unless you have a very fancy one

OpenStudy (anonymous):

so sqrt(6)i*(42)i Correct and using i^2 =-1 -> -sqrt(6)sqrt(42) and now u can just do the normal thing with square roots.